OpenStudy (anonymous):

Compute the directional derivative of the function f(x,y) = 4xy + 3y^2 at the point P(3,-4) in the direction of the vector <-3,-1>.

OpenStudy (sidsiddhartha):

fisrt find the gradient of\[f(x,y)=4xy+3y^2\]

OpenStudy (sidsiddhartha):

can u do it?

OpenStudy (anonymous):

No. It's new to me, can't figure out the right way to think of it.

OpenStudy (sidsiddhartha):

but you need to lnow how to find gradient, without this u cant solve it

OpenStudy (sidsiddhartha):

ok i can solve it for u ,do u know partial derivatives?

OpenStudy (anonymous):

Not really getting the hang of it. Is it this? x partial = 4y y partial = 4x + 6y

OpenStudy (sidsiddhartha):

yes so the gradient will be \[\nabla f(x)=4y \hat{i}+(4x+6y) \hat{j}\]

OpenStudy (anonymous):

so, plug in -3 and -1 for i and j?

OpenStudy (sidsiddhartha):

no now plug 3,-4 for x and y

OpenStudy (sidsiddhartha):

what u get ?

OpenStudy (sidsiddhartha):

\[\nabla f(x)=-14 \hat{i}+12 \hat{j}\]

OpenStudy (sidsiddhartha):

right?

OpenStudy (anonymous):

Yes. How do you enter the gradient symbol?

OpenStudy (sidsiddhartha):

just type "nabla" in the equation editor

OpenStudy (sidsiddhartha):

now the vector given is\[-3 \hat{i}- \hat{j}\] right?

OpenStudy (anonymous):

Yes.

OpenStudy (sidsiddhartha):

so now its unit vector will be--\[U_v=\frac{ -3 \hat{i}-\hat{j} }{ \sqrt{3^2+1^2} }=\frac{ -3 \hat{i}-\hat{j} }{ \sqrt{10} }\] right?

OpenStudy (sidsiddhartha):

ok?

OpenStudy (anonymous):

how?

OpenStudy (sidsiddhartha):

take a vector V=ai+jb then its unit vector is\[U_v=\frac{ the~vector }{ It's~magnitude }=\frac{ a \hat{i}+b \hat{j} }{ \sqrt{a^2+b^2} }\]

OpenStudy (anonymous):

ok

OpenStudy (sidsiddhartha):

now finnaly directional derivative is\[\large D.D=unit~vector*gradient_{at~the~given~point}\\=\frac{ -3 \hat{i}-\hat{j} }{ \sqrt{10} }*(-14 \hat{i}-12 \hat{j})=\frac{ 42+12 }{ \sqrt{10} }=\frac{ 54 }{ \sqrt{10} }\]

OpenStudy (sidsiddhartha):

ok?

OpenStudy (anonymous):

hmm.. Answer is not correct

OpenStudy (sidsiddhartha):

hey ahh we did a small mistake

OpenStudy (anonymous):

12 was supposed to be positive?

OpenStudy (sidsiddhartha):

\[[\nabla f(x)]_{(3,-4)}=-16 \hat{i}-12 \hat{j}\]

OpenStudy (sidsiddhartha):

whats the answer \[60/\sqrt{10}\]

OpenStudy (anonymous):

Ah. Yes, that's right. Thank you.

OpenStudy (sidsiddhartha):

yw!!